The subjects of a study by Dugoff et al. (A-5) were 10 obstetrics and gynecology interns at the
University of Colorado Health Sciences Center. The researchers wanted to assess competence in
performing clinical breast examinations. One of the baseline measurements was the number of such
examinations performed. The following data give the number of breast examinations performed for
this sample of 10 interns.
Intern Number No. of Breast Exams Performed
1 30
2 40
3 8
4 20
5 26
6 35
7 35
8 20
9 25
10 20
1
Expert's answer
2020-04-15T12:35:04-0400

"\bar x=25.9,s=9.4687, n=10."

"95\%CI=(\bar x-t_{0.025,n-1}\frac{s}{\sqrt{n}},\bar x+t_{0.025,n-1}\frac{s}{\sqrt{n}})="

"=(25.9-2.2622\frac{9.4687}{\sqrt{10}},25.9+2.2622\frac{9.4687}{\sqrt{10}})=(19.1264,32.6736)".